Daniel has to decide whether to sell lemonade or coffee. He's calculated that if he sells lemonade there is a 20% chance of getting $100 and an 80% chance of losing $20. If he sells coffee, there is a 40% chance of getting $80 and a 60% chance of losing $5. Investment costs are already included. Which should he sell?
Answer by stanbon(75887)

Daniel has to decide whether to sell lemonade or coffee.
He's calculated that if he sells lemonade there is a 20% chance of getting $100 and an 80% chance of losing $20.
Expected value for lemonade = 0.20*100 - 0.8*20 = 20-16 = $4.00
--------------------
If he sells coffee, there is a 40% chance of getting $80 and a 60% chance of losing $5. Investment costs are already included.
Expected value for coffee = 0.40*80 - 0.60*5 = 32 - 3 = $29.00
---------------------
Which should he sell?
Ans: Coffee
-------------
Cheers,
Stan H.
------------